Solve for x : 5/6 = 8/x
Answer by jim_thompson5910(35256)

5/6 = 8/x
5*x = 6*8 ... Cross multiply
5*x = 48
5*x/5 = 48/5 ... Divide both sides by 5
x = 48/5 ... Answer as a fraction
x = 9.6 ... Answer in decimal form